# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v1

@@ 6 The Impossibility Theorem
7 Escaping Impossibility
Understanding the limitations of the theoretical framework is crucial both for assessing the theorem’s scope and identifying potential paths for circumventing the impossibility result.
The Impossibility Theorem depends critically on the application of the Green-Laffont theorem and violation of the Semantic Information Conservation Principle. We should also explicitly establish that the LLM inference setting does indeed satisfy all necessary conditions required.
7.1 Applicability of Green-Laffont Theorem
The Green-Laffont theorem requires that agents have independently distributed private information. That is represented by the knowledge profiles $\theta=(\theta_{1},\ldots,\theta_{n})$ where:
$P(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i})$
This condition is satisfied in LLMs distributing knowledge (or its representation) across different components. Empirical evidence from mechanistic interpretability research (Meng et al. meng2023locatingeditingfactualassociations , Geva et al.geva2023dissecting ) demonstrates that different attention heads and MLP components specialize in distinct knowledge domains, supporting the independence assumption. However, it is also clear that transformer attention heads share residual streams, layer normalizations, and optimizer states, thereby creating strong correlations.
Surprisingly, as already stated, correlated knowledge assumption actually strengthens the impossibility result—correlated knowledge would amplify competition effects, making the impossibility more explicit. Milgorm and Weber milgrom1982theory showed that with their linkage principle. Roughgarden et al. roughgarden2016optimal provides arguments amplifying the violation of information conservation principle with interdependent knowledge.
The theorem also requires utilities to have the quasilinear form $u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta)$. In transformer-based LLMs, this corresponds to $v_{i}$ measuring contribution of component $i$ to reducing the model’s loss function, and $p_{i}$ accumulating computational/optimization cost imposed on that component. Real transformer training involves complex gradient interactions, and emergent behaviors that may resist simple decomposition. How that may affect the impossibility result remains to be verified.
The knowledge-constrained optimality property requires that the outcome $g(s^{*}(\theta))$ minimizes hallucination cost among feasible responses. That corresponds to the idea of optimization of aggregated preferences of competing agents under truthful knowledge representation.
Finally, the Green-Laffont theorem assumes that transfer payments between agents are feasible. In our setting that corresponds to payment function $p_{i}(s,\theta)$ representing transfers of semantic information among LLM components.
Given that all conditions are satisfied, the Green-Laffont theorem implies that any mechanism satisfying truthfulness and knowledge-constrained optimality must have a payment structure of the form:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})$ (31)
where $h_{i}(s_{-i})$ is some function independent of agent $i$’s strategy.
7.2 Applicability to LLM Inference
While actual LLM generation involves sequential token prediction with evolving context, repetitive generation of response to the same query justifies modelling inference as a sealed-bid auction. Also, the revenue equivalence theorems in auction theory suggest the conclusions we draw extend to other auction formats, including dynamic auctions more closely resembling autoregressive generation. Having said that, we must acknowledge all that demands further investigations.
There are many ways we can map agents to LLM architecture. One approach takes each attention head in a transformer as an agent with private knowledge encoded in the query, key, value, and output projection matrices ($W_{Q}$, $W_{K}$, $W_{V}$, $W_{O}$). Studies in mechanistic interpretability support that mapping. Circuit analysis (Elhage et al. elhage2021mathematical ) demonstrated that specific attention heads may specialize in distinct linguistic functions. For instance, some heads track subject-verb agreement, others detect named entities, and still others focus on coreference resolution. This specialization corresponds to our notion of agents with private knowledge.
The feed-forward networks in transformers can be linked to memory systems (Geva et al. geva2023dissecting ), storing factual associations. These components encode distributional knowledge about the world, activate specific neurons for particular factual domains, contribute to response generation through their activation patterns.
Recent works lindsey2025biology has also identified functional circuits spanning multiple layers and components that collectively implement specific capabilities. These circuits span across attention heads and MLP layers, perform specific reasoning tasks and activate as coherent units during inference.
7.3 Pathways to Circumvent Impossibility
Approximate Solutions: While perfect satisfaction of all four properties is impossible, approximate solutions might achieve near-optimal performance across all dimensions. The question becomes: how close to the theoretical limits can practical systems approach?
External Knowledge Integration: Our framework analyzes closed systems with fixed knowledge representations. Interactive systems incorporating real-time retrieval, human feedback, or external verification might alter the fundamental constraints by changing available knowledge states during inference.
Multi-Stage Architectures: Complex pipelines involving separate reasoning, verification, and generation stages could potentially distribute the trade-offs across different components, though each stage would still face internal impossibility constraints.
Relaxing Independence: If neural components can achieve perfect knowledge coordination (violating our independence assumption), the Green-Laffont theorem constraints might not apply. This suggests investigating architectures with explicit coordination mechanisms between attention heads and MLP blocks.
Compositional Reasoning: If models can perfectly decompose complex queries into independent sub-problems solvable without trade-offs, the impossibility might not apply to the overall system, even if it constrains individual components.
External Alignment: Human-in-the-loop systems, constitutional AI, or robust feedback mechanisms might provide external constraints that resolve the internal trade-offs, though this shifts rather than eliminates the fundamental tensions.
This completes the formal demonstration that perfect hallucination control faces fundamental mathematical limitations, regardless of architectural sophistication or computational resources.

### v2

@@ 6.1 Inference with Independent Private Knowledge
Theorem 2 describes LLM inference as a marketplace where knowledge is currency. The information contributions measures each agent’s market power, showing how much worse off others would be without agent $i$’s knowledge. When all agents have valuable knowledge, everyone contributes and that inevitably violates conservation of information. Then, new information must become accessible, but it may have no grounding in available knowledge. It may become lucky hallucination or knowledge-unconstrained imagination.
6.2 Inference with Probabilistic Predictions
To show impossibility in LLM transformer setting, where auction of ideas involves bids representing probability distributions, we refer to the theory of proper scoring developed by Savage savage1971elicitation , Gneiting and Raftery gneiting2007strictly , among others.
Definition 12 (Strictly Proper and Concave Scoring Rule).
Let $\mathcal{Y}$ be a finite or countably infinite outcome set (representing responses or intermediate conclusions). A scoring rule is a function $S:\Delta(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ that assigns a number $S(\pi,y)$ to probability distribution $\pi\in\Delta(\mathcal{Y})$ predicting random outcome $y\in\mathcal{Y}$.
We say the scoring rule $S$ is proper if:
$\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}\geq\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}$ (44)
for all distributions $\pi^{*},\pi\in\Delta(\mathcal{Y})$. The rule is strictly proper if equality holds only when $\pi=\pi^{*}$. We call the scoring rule concave, if $S(\tilde{\pi},y)$ is concave with respect to $\tilde{\pi}$ for every fixed $y\in\mathcal{Y}$.
In other words, scoring rule evaluates how good a prediction we can make of $y$ drawn from $\pi^{*}$ when we use distributions $\pi$. In particular, strictly proper scoring evaluates true distribution $\pi^{*}$ better (on average) than any other distribution $\pi$ we could use to predict the outcome.
The following result shows that we can use proper scoring to train LLM that reconstructs data distribution correctly.
Theorem 3 (Truthfullness under Strictly Proper Scoring Rules).
Let $S$ be a strictly proper scoring rule. Consider an agent encoding knowledge with $\pi^{*}\in\Delta(\mathcal{Y})$ and using probability distribution ${\pi}\in\Delta(\mathcal{Y})$ to generate prediction. If agent’s expected payoff from using ${\pi}$ is given by
$u(\pi|\pi^{*})=\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=\sum_{y\in\mathcal{Y}}\pi^{*}_{y}S(\pi,y),$ (45)
then truthful reporting ${\pi}=\pi^{*}$ is the unique optimal strategy.
Proof.
Because $S$ is proper, by definition $\pi^{*}$ maximizes the expected score. If $S$ is strictly proper, then for $\pi\neq\pi^{*}$ we have
$u(\pi^{*}|\pi^{*})=\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}>\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=u(\pi|\pi^{*}).$ (46)
Therefore, truthfully reporting $\pi^{*}$ is uniquely optimal. ∎
The scoring framework allows us to go beyond the important case of Green-Laffont/VCG model with independent private knowledge. With strictly-proper scores we address the general setting of probabilistic predictions that needs only occasional disagreement of probability distributions that agents use in the auction of ideas to formulate their thoughts (if we were to anthropomorphize for a moment). We will see in the following sections that transformer heads of modern LLMs implement predictive rules of that type and exhibit precisely such a disagreement. That makes the proper scoring framework well suited to explain the emergence of inference impossibilities in LLMs.
For that purpose, without loss of generality (and referring to the notion of superposition in mechanistic interpretability elhage_toy_2022 ), let us consider the following family of response generation mechanisms.
Definition 13 (Convex aggregation mechanism).
Given $H\geq 2$ agents encoding knowledge with probability distributions $\pi^{(i)}\in\Delta(\mathcal{Y})$ for $i=1,\dots,H$. The convex aggregation mechanism $\mathcal{M}$ generates response to a query based on the convex aggregate distribution:
$\Pi=\sum_{h=1}^{H}\beta_{h}\pi^{(h)},\ \text{where}\ \sum_{h=1}^{H}\beta_{h}=1\ \text{and}\ \beta_{h}>0.$ (47)
When outcome $y^{*}$ is generated, the agent’s information contribution is
$p_{h}=\beta_{h}\,[c-S(\pi^{(h)},y^{*})],$ (48)
where $c\geq\max_{h,y}S(\pi^{(h)},y)$ and $S$ is a proper scoring rule.
For that family of LLMs trained to optimize scoring, we can formulate the following general impossibility result.
Theorem 4 (Impossibility Theorem for Transformer).
For any query space $\mathcal{Q}$ containing non-trivial queries requiring at least two independent agents with mutually different predictive distributions, no LLM trained with strictly proper and concave scoring with convex aggregation of predictive distributions can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
Proof.
Each contribution $p_{h}=\beta_{h}[c-S(\pi^{(h)},y^{*})]$ is a positive-affine transformation of the strictly proper score $S$. Since scaling by positive constant $\beta_{h}>0$ and adding constant $c$ preserves the arg-max, truthful reporting maximizes each agent’s expected utility by Theorem 3.
Because $S(\cdot,y^{*})$ is strictly concave and at least two beliefs differ, Jensen’s inequality gives:
$S\left(\Pi,y^{*}\right)=S\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (49)
Therefore, there exists a gap, following from Jensen’s inequality for concave functions, equal to:
$\Gamma=S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})>0.$ (50)
Expanding the sum of all payments, we get:
$\sum_{h=1}^{H}p_{h}$ $=\sum_{h=1}^{H}\beta_{h}[c-S(\pi^{(h)},y^{*})]$ (51)
$=c\sum_{h=1}^{H}\beta_{h}-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})=c-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (52)
However, adding and subtracting $S(\Pi,y^{*})$ shows that is equivalent to:
$\sum_{h=1}^{H}p_{h}$ $=[c-S(\Pi,y^{*})]+[S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})]=[c-S(\Pi,y^{*})]+\Gamma.$ (53)
Since $c\geq\max_{h,y}S(\pi^{(h)},y)\geq S(\Pi,y^{*})$ and $\Gamma>0$, we conclude that:
$\sum_{h=1}^{H}p_{h}=[c-S(\Pi,y^{*})]+\Gamma>0$ (54)
Information conservation demands $\sum_{h=1}^{H}p_{h}=0$, but we have just shown $\sum_{h=1}^{H}p_{h}>0$ whenever beliefs differ and the scoring rule is strictly concave. Therefore, the four properties cannot coexist. ∎
This theorem provides the foundation for extending impossibility results to probabilistic settings. The key insight is that proper scoring rules drive truth-revealing responses which promote using true (trained knowledge based) beliefs in the inference process.
That explains why cross-entropy loss (negative log-score) is so effective in training neural networks. It creates gradients that push each component toward truthful probability reporting. Every attention head in a transformer is implicitly playing this truth-telling game.
However, while that ensures truthfulness and relevant knowledge revelation, it doesn’t prevent collective hallucination. Each head truthfully reports its partial view, but their aggregation can still create false certainty.
7 The Emergence of Impossibility in Transformer Architectures
To demonstrate where the impossibility may emerge in transformer architectures, let us analyze a minimal yet representative example of a tiny transformer predicting the next token in the famous English pangram:
The quick brown fox jumps over the lazy dog.
We trace the inference process, inspect every mathematical operation from input embeddings through final predictions, showing exactly where the impossibility emerges.
7.1 Model Architecture and Input Representation
Our micro-transformer has the following specifications:
Parameter Value Description
$d_{m}$ 4 Model dimension
$H$ 2 Number of attention heads
$d_{k}$ 2 Query/Key dimension per head
$d_{v}$ 2 Value dimension per head
$d_{f\!f}$ 6 Feedforward hidden dimension
$|V|$ 6 Vocabulary size
$T$ 3 Sequence length
Table 1: Micro-transformer architecture parameters
For the sake of simplicity, consider the following vocabulary:
$V=\{\langle\text{PAD}\rangle,{The},{quick},{brown},{fox},{dog}\},$
including padding token (with index zero in $V$). Each token (representing word for simplicity) in that vocabulary is represented by a row in the embedding matrix
$\mathbf{E}=\begin{bmatrix}0&0&0&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}.$
The un-embedding matrix is defined as $\mathbf{U}=\mathbf{E}^{\top}$.
Consider the following input sequence
The quick brown
It is represented by the input matix:
$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix},$
where each row corresponds to the embedding of tokens at positions 1, 2, and 3.
The following system of equations defines output $\mathbf{o}_{t}^{(h)}$ of attention head $h$ at position $t$:
[…]
The block per-token multihead output $\mathbf{a}_{t}$ is then given by the sum:
[…]
The attention logits, i.e., the scores whose exponentials become probabilities, are given by:
$\mathbf{l}^{(h)}=\mathbf{o}^{(h)}\mathbf{U},\quad\text{so that}\quad\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}.$ (57)
Next, we take into account the feedforward network (FFN) introducing nonlinear correction:
$\mathbf{z}=\mathbf{W}_{2}\max(0,(\mathbf{X}_{0}+\mathbf{a})\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (58)
generating the FFN logits:
$\mathbf{L}_{f}=\mathbf{z}\mathbf{U}.$ (59)
As a result, the residual stream is adjusted as follows:
$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{a}+\mathbf{z}.$ (60)
Finally, the probability distribution $\Pi$ of the next token is calculated by aggregating attention and FFN logits into:
$\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}\quad\text{and normalizing into}\quad\Pi=\text{softmax}(\mathbf{L}).$ (61)
7.2 The Inference Process
We now analyze the inference process step by step to understand the moments the impossibility result may emerge. We shall keep things simple but realistic, recreating the essential steps of the inference process in transformer block.
Step 1: Query and Key
There are $H=2$ heads for which we define the query and key, or better yet, demand and supply bids in the auction of ideas. Query announces demand for information (what does this token need?), key submits supply offers (what does the previous tokens offer!). Then those bids are compared to find good matching (and clear the information market).
Let us assume the current position is $t=3$, so we attend to brown. Then, consider the following knowledge encoded in each head:
$\mathbf{W}_{Q}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&1\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(1)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix},$ (62)
$\mathbf{W}_{Q}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(2)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}.$ (63)
Matrix $\mathbf{W}_{Q}^{(h)}$ amplifies the coordinates of tokens in $\mathbf{X}$ that contribute to relevant context. As a result, we get a query vector that represents the following question:
what information is token $\mathbf{X}_{t}$ looking for?
In other words, the query vector indicates how strongly the current token demands information along the directions in $\mathbf{Q}_{t}^{(h)}$.
Then computing queries for $t=3$ with $\mathbf{X}_{3}$ yields:
$\mathbf{Q}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{Q}_{3}^{(1)}=[1,0].$
Head 1 (agent) demands information stored along direction $[0,1]$. Similarly, Head 2 (agent) demands information stored along direction $[1,0]$. That means the heads develop different (orthogonal) privately known and independent attention patterns, leading to beliefs about the next token that are as different as possible.
Next, we need to see what information can be provided by the input vector given the knowledge encoded in $\mathbf{W}_{K}^{(h)}$. Given that knowledge, we get from each head (agent) the key vectors that inform:
along which directions do the available tokens in $\mathbf{X}_{\leq t}$ provide information relevant for token $\mathbf{X}_{t}$.
Since we have assumed, for simplicity, that $\mathbf{W}_{Q}^{(h)}=\mathbf{W}_{K}^{(h)}$, the query and key vectors are perfectly aligned:
$\mathbf{K}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{K}_{3}^{(1)}=[1,0]\quad\text{and}\quad\mathbf{K}_{<3}^{(h)}=[0,0].$
Another interesting way to look at query and key is this. It is a memory lookup mechanism. Keys are address in memory, query selects addresses in memory which probably store relevant content. As we will see next, that content is provided by the value vectors.
Step 2: Attention Score and Weights
Given the demand and supply bids providing directions along which information is required and stored, in the next step the attention mechanism executes the bid matching procedure. Namely, the attention scores are computed to measure cosine similarity of bids:
$\text{scores}_{3}^{(h)}=\frac{\mathbf{Q}_{3}^{(h)}(\mathbf{K}_{\leq 3}^{(h)})^{\top}}{\sqrt{d_{k}}}.$ (64)
In our specific example, we have:
$\text{scores}_{3}^{(1)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707]\quad\text{and}\quad\text{scores}_{3}^{(2)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707].$
Applying softmax to get attention weights:
$\boldsymbol{\alpha}_{3}^{(h)}=\text{softmax}(\text{scores}_{3}^{(h)}),$ (65)
we get the following probability distributions (weights) over relevant context information:
$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50].$
Therefore, based on the encoded knowledge, heads attend strongly to position 3 (brown).
Step 3: Context Vector Computation
The context vector for each head is calculated as a linear combination ($\boldsymbol{\alpha}$ is a simplex) of value vectors:
$\mathbf{h}_{3}^{(h)}=(\boldsymbol{\alpha}_{3}^{(h)})^{\top}\mathbf{V}_{\leq 3}^{(h)}.$ (66)
The value vectors,
$\mathbf{V}_{\leq t}^{(h)}=\mathbf{X}_{\leq t}^{(h)}\mathbf{W}^{(h)}_{V},$ (67)
prepare context information relevant for the next token prediction (performed in next steps). That critical content is extracted from the input vector based on the trained knowledge stored in $\mathbf{W}^{(h)}_{V}$. Columns in $\mathbf{W}^{(h)}_{V}$ detect presence of relevant information in $\mathbf{X}_{\leq t}$ and scale it appropriately.
We have:
$\mathbf{V}_{\leq 3}^{(1)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\end{bmatrix}\ \text{and}\ \mathbf{V}_{\leq 3}^{(2)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\end{bmatrix}.$
That means, both heads believe that the presence of brown in position 3 is relevant (row 3 in $\mathbf{W}^{(h)}_{V}$ has nonzero elements). Token brown statistically predicts important pattern extracted from training data. That information is encoded by both heads as a vectors in value space,
$\mathbf{V}_{3}^{(1)}=[0,3]\quad\text{and}\quad\mathbf{V}_{3}^{(2)}=[1.2,0].$
Given the demand-supply matching encoded in $\boldsymbol{\alpha}^{(h)}_{3}$, that yields head context vectors:
$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$
Step 4: Output Projection and Feedforward Exploration
Each head then projects its context vector to residual space based on the privately held information routing knowledge encoded in $\mathbf{W}_{O}^{(h)}$. That private knowledge translates (or transmits) valuable context information into embedding vectors. Namely, each head calculates:
$\mathbf{o}_{3}^{(h)}=\mathbf{h}_{3}^{(h)}\mathbf{W}_{O}^{(h)}.$ (68)
With trained matrices:
$\mathbf{W}_{O}^{(1)}=\begin{bmatrix}0&0&0&0\\ 0&0&1&0\end{bmatrix}\quad\text{and}\quad\mathbf{W}_{O}^{(2)}=\begin{bmatrix}0&0&0&1\\ 0&0&0&0\end{bmatrix},$ (69)
the attention mechanism generates ranking of evidence emphasizing relevance of the vocabulary tokes:
$\mathbf{o}_{3}^{(1)}=[0,0,0,1.5,0]\quad\text{and}\quad\mathbf{o}_{3}^{(2)}=[0,0,0,0,0.6].$
So, in the next token prediction auction of ideas, Head 1 votes for fox and Head 2 votes for dog.
The votes are added into residual dimensions to adjust the meaning of input embedding. For that, the outputs (votes) are aggregated into:
$\mathbf{a}_{3}=\sum_{h=1}^{H}\mathbf{o}_{3}^{(h)}=[0,0,0,1.5,0.6]\quad\text{and}\quad\mathbf{a}_{<3}=[0,0,0,0,0],$
and added to input vector:
$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{a}.$
Than completes operations of attention block.
The next step is to add contribution of FFN feedforward block (or another agent participating in the aution). We have:
$\mathbf{X}_{2}=\mathbf{X}_{1}+\mathbf{W}_{2}\max(0,\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2}.$
That FFN operation explores context locally and introduces additional innovations in meaning, shaping the next token prediction result. It can amplify or counteract the votes injected by attention. Notice that adjustments are applied at every position separately, so that ideas provided by attention bids are interpreted and rewritten by the FFN agent.
7.2.1 Step 5: Logits and Prediction
Finally, the unembedding matrix $\mathbf{U}$ maps the multi-headed attention and FFN outputs from model space to vocabulary in order to calculate probability distribution for the next token prediction.
Each head’s contribution to the final logits is:
$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6].$
The total attention logits are the sum of individual head logits:
$\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}=[0,0,0,0,1.5,0.6].$
That sum is combined with FFN logits $\mathbf{L}_{f}$ that introduce some creative noise. As a result, we obtain:
$\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52].$
That vector translates into the final probability distribution:
$\Pi=\text{softmax}(\mathbf{L})=[0.1,0.1,0.13,0.12,\mathbf{0.38},0.17].$
The winner of the auction of ideas is selected based on $\Pi$. The inference process generates and collects competing beliefs about relevant context and next token prediction, all based on the knowledge stored in the attention heads and FFN layer. And so, we see that when that knowledge is used, the most probable next token (word) reachable in the vocabulary is
$\text{argmax}\ \Pi=\textit{fox}.$
7.3 The Impossibility Theorem for Transformer
With the deep insight into the transformer inference process, we are ready to show the impossibility in that special and important setting.
Theorem 5 (Impossibility for Transformers).
In transformer LLMs with non-redundant attention heads, it is impossible to implement inference process in which each head outputs its true conditional distribution while conserving semantic information and minimize aggregate loss with logit summation.
Proof.
Let attention heads $h=1,\ldots,H$ output logits $\mathbf{l}^{(h)}\in\mathbb{R}^{|\mathcal{V}|}$. Then, individual head distributions are given by:
$\pi^{(h)}_{y}=\frac{\exp\bigl{(}l^{(h)}_{y}\bigr{)}}{Z_{h}}\ \text{with}\ Z_{h}=\sum_{\bar{y}}\exp\bigl{(}l^{(h)}_{\bar{y}}\bigr{)}.$ (70)
Similarly, the aggregate distribution from summed logits is calculated as:
$\Pi_{y}=\frac{\exp(L_{y})}{Z}\ \text{with}\ Z=\sum_{\bar{y}}\exp(L_{\bar{y}})\ \text{and}\ L_{y}=\sum_{h=1}^{H}l^{(h)}_{y}.$ (71)
For any observed (realized) token $y^{*}$, the head losses and aggregate loss are:
$p_{h}=-\log\pi^{(h)}_{y^{*}}=-l^{(h)}_{y^{*}}+\log Z_{h}\quad\text{and}\quad p_{H+1}=-\log\Pi_{y^{*}}=-L_{y^{*}}+\log Z.$ (72)
But that implies,
$\sum_{h=1}^{H}p_{h}=p_{H+1}+\gamma,\ \text{where }\Gamma=\sum_{h=1}^{H}\log Z_{h}-\log Z.$ (73)
If heads produce non-proportional logit vectors, then $\Gamma>0$ by convexity of log-sum-exp aggregation.
Therefore,
$\sum_{h=1}^{H}p_{h}>0,$ (74)
which violates semantic information conservation principle.
This final theorem bridges earlier impossibility results to concrete transformer architectures. It shows that concavity of scoring functions is the mathematical source of hallucination. The Jensen gap $\Gamma$ becomes a measurable quantity showing that probability aggregation violates conservation and any sufficiently capable transformer-based LLM must violate conservation of per-token information on non-trivial queries.
Let us note that the aggregate distribution cannot create Shannon information ex nihilo. Instead it reorganizes existing information to be more useful for the specific task of predicting $y^{*}$. Given fixed head logits, the Shannon entropy of a function (convex aggregate) of those logits is bounded from above by the the Shannon entropy of the fixed logits. This way information becomes more accessible, not more existent, as postulated by the reasoning with emergence operator and Theorem 1. The Jensen’s gap measures point-wise excess confidence of the predicted outcome, possibly making LLMs appear to know more (when in fact they do not).
8 Summary
This paper establishes a fundamental mathematical impossibility at the heart of large language model inference: no LLM can simultaneously achieve perfect hallucination control while maintaining semantic coherence and optimal knowledge utilization. That result demonstrates that four essential properties of ideal language model behavior are mutually incompatible, creating inevitable trade-offs that no architectural innovation or training methodology (aligned with the impossibility theorem assumptions) can eliminate.
For any query space $\mathcal{Q}$ containing non-trivial queries requiring integration of distributed knowledge, no language model inference mechanism can simultaneously:
  • represent encoded knowledge accurately,
  • preserve total information content during inference,
  • reveal all accessible relevant information,
  • minimize hallucination cost within available knowledge.
This impossibility emerges from the fundamental mathematical structure of information aggregation, not from any particular architectural choice.
8.1 Mathematical Foundation and Proof Strategy
The proof operates across three complementary mathematical frameworks, each capturing different aspects of LLM inference.
Auction-Theoretic Framework
We model LLM inference as an auction of ideas where neural components (attention heads, circuits, activation patterns) compete to contribute knowledge to response generation. Under the assumption of independent private knowledge profiles, we apply the Green-Laffont theorem to prove impossibility.
Probabilistic Framework
We extend the impossibility to any system where neural components output probability distributions by leveraging Savage’s theory of proper scoring rules. For convex aggregation mechanisms with strictly concave scoring functions, Jensen’s inequality creates an inevitable information aggregation gap. The Jensen’s gap represents excess confidence generated by probabilistic aggregation, violating information conservation whenever beliefs differ across components.
Transformer Architecture
We demonstrate how the impossibility manifests in actual transformer implementations through log-sum-exp convexity. The (Jensen’s) convexity gap $\Gamma$ provides a directly measurable quantity in transformer training, showing that information conservation violations are not merely theoretical but empirically observable.
8.2 Key Mathematical Contributions
To establish the impossibility result, we have designed a mathematical framework providing both formal and philosophical ground for the study. The Semantic Information Measure is a context-dependent information metric that captures bounded reasoning capabilities. The Emergence Operator models how reasoning makes hidden knowledge accessible within computational constraints. And Auction-Theoretic Modeling allows us to identify logical limitations of the inference process.
The impossibility results apply across multiple mathematical models of knowledge representations, training goals, and neural architectures. This universality suggests the trade-offs we have identified are fundamental features of distributed intelligence systems, not artifacts of particular modeling choices.
8.3 Implications for AI Development
Despite its name, the impossibility theorem provides constructive guidance for LLM design and deployment. Rather than pursuing impossible simultaneous optimization, we may consider balancing competing objectives. Namely, prioritize truthfulness and information conservation in safety-critical applications, allow controlled information conservation violations to enable innovative responses in creative applications, develop adaptive mechanisms that adjust trade-offs based on query context in general-purpose systems.
The impossibility result may guide LLM evaluation. Impossibility-aware metrics can measure trade-off quality rather than absolute performance, alignment strategies can negotiate trade-offs with human preferences rather than assuming universal optimality.
8.4 Significance for AI Theory
Just as Goedel’s incompleteness theorems revealed fundamental limits in formal systems, the impossibility theorems in this paper identify inherent constraints in knowledge integration and reasoning. Hallucination, or imagination, is a price of reasoning, much like incompleteness is a price of consistency.
The Jensen’s gap controlling the impossibility strength may provide insight into the relationship between aggregation and information content. The apparent creativity in large language models may be better understood as necessary violations of information conservation rather than genuine knowledge generation.
8.5 Limitations and Future Directions
The impossibility results apply specifically to non-trivial queries requiring knowledge integration from multiple sources, finite computational budgets that create bounded reasoning constraints, and distributed knowledge representations where information is spread across multiple components. Future work should investigate whether alternative knowledge architectures might circumvent some aspects of the impossibility.
8.6 Conclusion
The fundamental impossibility of perfect hallucination control represents both a limitation and a insight for the field of artificial intelligence. By understanding that trade-offs are mathematically inevitable, we can move beyond attempts at simultaneous optimization toward principled management of inherent tensions in intelligent systems. Formal definitions of the impossibility building blocks open new avenues for research at the intersection of mathematical analysis, engineering design, and philosophical inquiry into the nature of machine intelligence.
When we observe LLMs appearing to generate novel insights, they are either $(i)$ revealing information already implicit in their training (making knowledge accessible) or $(ii)$ violating conservation (hallucinating). This explains why LLMs can seem both remarkably creative and fundamentally limited. They excel at revealing hidden connections but cannot easily transcend information content of their training dataset.
This impossibility is not a bug but a feature of distributed intelligence. That price of knowledge integration in reasoning is worth paying.